OpenStudy (anonymous):

What % of 105 is 16.8?

OpenStudy (a1234):

x/100 = 16.8/105 1680/105 = ?

OpenStudy (anonymous):

For any math problem I use Tiger Algebra Calculator it has help me with soo many of my math problems

OpenStudy (timaashorty):

\[16 percent \] Douple check. \[16 x 105 = 1680 \]